OpenStudy (anonymous):

HELP ME PLEASE SOMEONE!!! THIS IS URGENT!!! Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3. a. -185 b.-77 c.-23 d.85

OpenStudy (anonymous):

is 3x3 mean power of x is 3..???that is 3x^3..???is it..??

OpenStudy (anonymous):

yah sorry those are powers

OpenStudy (anonymous):

x^3-2y^2-3x^3+z^4

OpenStudy (anonymous):

please come back??!! D':

OpenStudy (anonymous):

yeahh m back.. just one minute.

OpenStudy (anonymous):

ans is option C...i.e. -23

OpenStudy (anonymous):

an you Find x - 3y if x = 3 and y = -2.

OpenStudy (anonymous):

its 3-3(-2)=3+6=9